A survey of 2,150 adults reported that 53% watch news videos. Complete parts (a) through (c) below. a. Suppose that yotakeaample of 50 adults. If the population proportion of adults who watch news videos is 0.53, what is the probability that fewer than half in your sample will watch news videos? The probability is that fewer than half of the adults in the sample will watch news videos. (Round to four decimal places as needed.) b. Suppose that you take a sample of 250 adults. If the population proportion of adults who watch news videos is 0.53, what is the probability that fewer than half in your sample will watch news videos? The probability ishat fewer than half of the adults in the sample will watch news videos. Round to four decimal places as needed.) c. Discuss the effect of sarmple size on the sampling distribution of the proportion in general and the effect on the probabilities in parts (a) and (b). Choose the correct answer below. 2 OA. Increasing the sample size by a factor of 5 increases the standard error by a factor of 5. This causes the sampling distribution of the proportion to become more concentrated around the true population proportion of 0.53 and increases B. Increasing the sample size by a factor of 5 decreases the standard error by a factor of 5. This causes the sampling distribution of the proportion to become more concentrated around the true population proportion of 0.53 and C. The probabilities in parts (a) and (b) are the same. Increasing the sample size does not change the sampling distribution of the proportion. 2 the probability in part (b). decreases the probability in part (b)
Answer #1

Given: p 0.53, q 1-p 0.47,n-50 Taking normal approximation to binomial, the sampling distribution of sample proportion follow0.3336 0.43 P(p< 0.5)-P(Z< -0.43) 0.3336 (from z-table) P(2-0.43): in a z-table having area to the left of z, locate -0.4 inwe convert this to standard normal using z- 0.5 (0.53) z = 0.031566 -0.950396 -0.95 0.95 P(ρ < 0.5) = P(Z <-0.95) = 0.1711 (f
Answer #2

A survey of 2,250 adults found that 56% got news both online and offline

during a typical day.

a. Suppose that you take a sample of 50 adults. If the population

proportion of adults who get news both online and offline in a typical

day is 0.56, what is the probability that fewer than half in your sample

will get news both online and offline in a typical day?

b. Suppose that you take a sample of 250 adults. If the population

proportion of adults who get news both online and offline in a typical

day is 0.56, what is the probability that fewer than half in your sample

will get news both online and offline in a typical day?

c. Discuss the effect of sample size on the sampling distribution of the

proportion in general and the effect on the probabilities in parts (a)

and (b).


answered by: anonymous